# Readout-Side Bypass for Residual Hybrid Quantum-Classical Models

### v1

@@ I Introduction
Quantum machine learning (QML) offers promising advantages, including compact representations in high-dimensional Hilbert spaces and intrinsic support for nonlinear feature extraction via parameterized quantum circuits (PQCs) [li2025quantum, khurana2024quantum, xu2025quantum, zaman2023survey]. However, practical QML faces a core limitation: the measurement bottleneck, where high-dimensional classical inputs are compressed into a small number of quantum observables. This severely constrains downstream accuracy [tomar2025comprehensive, rohe2025quantum].
To mitigate this, hybrid quantum–classical models have been proposed, wherein classical neural networks process measured quantum features [havlivcek2019supervised, liu2021hybrid, de2022survey]. Yet, most hybrids remain bottlenecked—only the quantum features $Q(x)$ are passed forward, discarding original inputs $x$  [schuld2020circuit]. This narrow interface not only reduces performance but also amplifies susceptibility to membership inference attacks (MIA), as compressed features tend to overfit and leak more information  [farokhi2024maximal].
In classical deep learning, residual connections help preserve input information across layers and stabilize training [gomez2017reversible, zhang2019fixup]. However, such residuals bypass internal layers—not the terminal bottleneck. In QML, the information loss occurs at the quantum measurement stage, and traditional residuals are insufficient to address it.
Prior efforts to expand Q(x)Q(x)—via deeper PQCs, multi-basis measurements, or learned readouts [campbell2024deeperpqc, tasar2025multibasis, chatterjee2025readout]—reshape the feature space but do not expose the original input to the classifier. Moreover, matched-parameter comparisons and privacy evaluations are rarely reported, making it unclear whether such methods are accurate, compact, and robust.
@@
To evaluate the downstream implications of this design, we analyze its privacy robustness under common threat models. We consider a model-release adversary with full access to trained weights, performing MIA via threshold or shadow-model attacks [geiping2020inverting]. Reconstruction performance is measured by AUC; empirical robustness corresponds to AUC $\approx 0.5$, indicating indistinguishability.
On four benchmarks (Wine [wine1992], Breast Cancer [breastcancer1995], Fashion-MNIST Subset [xiao2017fashion], and Forest CoverType Subset [blackard1999comparative]), our residual hybrid achieves near-classical accuracy with 10–20% fewer parameters and significantly better privacy robustness (MIA AUC $\approx$ 0.5), outperforming both quantum-only and standard hybrid baselines. Ablation confirms that exposing both $x$ and $Q(x)$ to the classifier is crucial for overcoming the measurement bottleneck.
@@ III-A Experimental Setup
Datasets.  We evaluate on four datasets with diverse size and dimensionality: (A) Wine [wine1992]: 178 samples, 13 features, 3 classes; (B) Breast Cancer [breastcancer1995]: 569 samples, PCA-reduced from 30 to 10, binary; (C) Fashion-MNIST Subset [xiao2017fashion]: 3,000 samples, PCA-reduced to 16D, 3 classes; (D) Forest CoverType Subset [blackard1999comparative]: 5,000 samples, 12 features, 3 classes. All features are scaled to $[0,1]$. We apply 5-fold cross-validation in centralized settings and 80/20 splits per client in federated settings.
@@ III-B Baselines and Metrics
Baselines.  We compare with four baselines: (1) Federated Averaging (FedAvg) [mcmahan2017communication], the standard classical FL method; (2) Centralized training, where all data is pooled for global model training; (3) Local-only training, with no model sharing between clients; (4) Differentially Private FL (DP-FL) using Gaussian noise ($\epsilon=1.0$, $\delta=10^{-5}$). All models use the same classifier architecture and are matched in parameter count for fair comparison.
@@
  • Gradient inversion attacks [geiping2020inverting]: reconstruction quality is measured by PSNR; lower PSNR means stronger protection.
@@ V Conclusion
Our design requires no changes to the quantum circuit or encoding strategy, making it protocol-agnostic and easily pluggable into existing hybrid quantum-classical systems. The resulting models show promising improvements in convergence speed and privacy robustness, suggesting that strategic architectural changes - even at the readout interface - can have a meaningful impact on QML performance and security. Future work includes applying this method to larger models, real quantum hardware, and broader security contexts such as adversarial robustness and differential privacy.

### v3

@@ I Introduction
Quantum machine learning (QML) offers promising advantages, including compact representations in high-dimensional Hilbert spaces and intrinsic support for nonlinear feature extraction via parameterized quantum circuits (PQCs) [11, 10, 20, 21]. However, practical QML faces a core limitation: the measurement bottleneck, where high-dimensional classical inputs are compressed into a small number of quantum observables. This severely constrains downstream accuracy [17, 14].
To mitigate this, hybrid quantum–classical models have been proposed, wherein classical neural networks process measured quantum features [9, 12, 5]. Yet, most hybrids remain bottlenecked—only the quantum features $Q(x)$ are passed forward, discarding original inputs $x$  [15]. This narrow interface not only reduces performance but also amplifies susceptibility to membership inference attacks (MIA), as compressed features tend to overfit and leak more information  [6].
In classical deep learning, residual connections help preserve input information across layers and stabilize training [8, 22]. However, such residuals bypass internal layers—not the terminal bottleneck. In QML, the information loss occurs at the quantum measurement stage, and traditional residuals are insufficient to address it.
Prior efforts to expand Q(x)Q(x)—via deeper PQCs, multi-basis measurements, or learned readouts [3, 16, 4]—reshape the feature space but do not expose the original input to the classifier. Moreover, matched-parameter comparisons and privacy evaluations are rarely reported, making it unclear whether such methods are accurate, compact, and robust.
@@
To evaluate the downstream implications of this design, we analyze its privacy robustness under common threat models. We consider a model-release adversary with full access to trained weights, performing MIA via threshold or shadow-model attacks [7]. Reconstruction performance is measured by AUC; empirical robustness corresponds to AUC $\approx 0.5$, indicating indistinguishability.
On four benchmarks (Wine [1], Breast Cancer [18], Fashion-MNIST Subset [19], and Forest CoverType Subset [2]), our residual hybrid achieves near-classical accuracy with 10–20% fewer parameters and significantly better privacy robustness (MIA AUC $\approx$ 0.5), outperforming both quantum-only and standard hybrid baselines. Ablation confirms that exposing both $x$ and $Q(x)$ to the classifier is crucial for overcoming the measurement bottleneck.
@@ III-A Experimental Setup
Datasets.  We evaluate on four datasets with diverse size and dimensionality: (A) Wine [1]: 178 samples, 13 features, 3 classes; (B) Breast Cancer [18]: 569 samples, PCA-reduced from 30 to 10, binary; (C) Fashion-MNIST Subset [19]: 3,000 samples, PCA-reduced to 16D, 3 classes; (D) Forest CoverType Subset [2]: 5,000 samples, 12 features, 3 classes. All features are scaled to $[0,1]$. We apply 5-fold cross-validation in centralized settings and 80/20 splits per client in federated settings.
@@ III-B Baselines and Metrics
Baselines.  We compare with four baselines: (1) Federated Averaging (FedAvg) [13], the standard classical FL method; (2) Centralized training, where all data is pooled for global model training; (3) Local-only training, with no model sharing between clients; (4) Differentially Private FL (DP-FL) using Gaussian noise ($\epsilon=1.0$, $\delta=10^{-5}$). All models use the same classifier architecture and are matched in parameter count for fair comparison.
@@
  • Gradient inversion attacks [7]: reconstruction quality is measured by PSNR; lower PSNR means stronger protection.
@@ V Conclusion
Our design requires no changes to the quantum circuit or encoding strategy, making it protocol-agnostic and easily pluggable into existing hybrid quantum-classical systems. Future work includes scaling to real quantum hardware and broader security contexts.